# Multi-Physics-Enhanced Bayesian Inverse Analysis: Information Gain from Additional Fields

### v1

@@ 2.1 Coupling in multi-physics models
Besides a multi-physics log-likelihood function, we also need a computational model $\mathcal{M}$ that models each of the observed physical fields. Such a multi-physics model must account for the coupling between the fields. Regarding inverse analysis, the type of coupling influences the information about the uncertain parameters that observations of each physical field can provide. In this section, we take a closer look at the coupling types in multi-physics models with two fields and their potential to improve an inverse analysis.
@@
The third coupling type that provides inverse information about the parameters $\bm{x}$ in the second field $\bm{y}_{2}$ is presented in the right diagram of Figure 4. Here, both output fields $\bm{y}_{1}$ and $\bm{y}_{2}$ depend on each other, thus the name fully coupled. An explicit dependency of the second field $\bm{y}_{2}$ on the parameters $\bm{x}$ is optional. Implementing and evaluating fully coupled models is challenging due to the larger system matrices and interdependencies between fields. Depending on the coupling strength, these models often require advanced numerical solution strategies to ensure stability and efficiency (see, e.g., [32]).
Uncoupled One-Way Coupled Fully Coupled
Elastic deformation:
$\bm{x}$:1 Young’s modulus
$\bm{y}_{1}$: Deflection under force
$\bm{y}_{2}$: Vibroacoustic signal
Fluid dynamics with smoke:
$\bm{x}$:1 Fluid viscosity
$\bm{y}_{1}$: Fluid velocity
$\bm{y}_{2}$: Smoke density
Fluid-structure interaction:
$\bm{x}$:1 Fluid parameters
$\bm{y}_{1}$: Fluid velocity
$\bm{y}_{2}$: Structure deformation
Medical imaging:
$\bm{x}$:1 Density field
$\bm{y}_{1}$: Ultrasound imaging
$\bm{y}_{2}$: Computed tomography (CT) scan
Heart contraction [18]:
$\bm{x}$:1 Electrical state variables
$\bm{y}_{1}$: Electrocardiogram
$\bm{y}_{2}$: Myocardium displacement
Natural convection:
$\bm{x}$:1 Thermal parameters
$\bm{y}_{1}$: Temperature field
$\bm{y}_{2}$: Fluid velocity
Marine electromagnetics [12]:
$\bm{x}$:1 Seawater & subsurface resistivities
$\bm{y}_{1}$: Magnetotellurics
$\bm{y}_{2}$: Controlled source electromagnetics
Lung ventilation [19]:
$\bm{x}$:1 Lung stiffness field
$\bm{y}_{1}$: Air flow at ventilator
$\bm{y}_{2}$: EIT voltages
Thermo-structure interaction [32]:
$\bm{x}$:1 Thermal stress field
$\bm{y}_{1}$: Structure deformation
$\bm{y}_{2}$: Temperature field
Information about the coupling type of a multi-physics model is provided by the model’s system matrix. We will briefly address this for multi-physics models with two output fields, although the following concepts also apply to models with an arbitrary number of physical fields [32]. The physical fields of interest are typically governed by nonlinear partial differential equations that cannot be solved analytically. Instead, their solutions are approximated at discrete points in space and time. After discretization, a fully coupled model with two physical fields can be formulated as a set of nonlinear algebraic equations:
@@ 3 Information gain in multi-physics-enhanced Bayesian inverse problems
In this work, we measure the uncertainty reduction from the prior to the posterior distribution using the information gain. This measure quantifies the amount of information gained about the parameters $\bm{x}$ by including the observations $\bm{y}_{\text{obs}}$. The information gain is a well-established measure in information theory. It is also widely used in Bayesian optimal experimental design to find an experimental design that maximizes the expected information gain [33, 2, 34, 35, 36]. In contrast to this use case, we use the information gain here to assess the additional reduction in uncertainty in a Bayesian inverse problem achieved by adding observations of a second physical field.
@@
% Introduce entropy and information gain The information gain is based on the difference in (Shannon) entropy between the prior and the posterior distributions. The entropy \$H(x)\$ quantifies the expected information \$I(x)\$ in a random variable \$x\$ with probability distribution \$p(x)\$ [37]:
@@
This work relies on entropy instead of variance to measure uncertainty. While the variance measures only the spread of a random variable around its mean, the entropy captures the uncertainty across the entire support of its probability distribution. % Moreover, the entropy reflects our intuition of uncertainty as the variance of a random variable with a single mode decreases. If a parameter is highly uncertain, we expect its uncertainty to decrease slowly with decreasing variance. However, if the uncertainty in the parameter is already low, we expect a stronger decrease in uncertainty as the variance decreases further. The entropy exhibits this behavior for a univariate, normally distributed random variable, as shown in Figure 5. % Based on these considerations, we rely on the entropy to compare the uncertainty in the prior and posterior distributions of an inverse problem. In particular, we consider the information gain \$IG(y_obs)\$, which is the relative entropy or Kullback-Leibler divergence between the prior distribution \$p(x)\$ and the posterior distribution \$p(x|y_obs)\$ [38]:
@@ 4 Numerical demonstration
% Goal of the numerical demonstration In this section, we demonstrate the potential of multi-physics observations and modeling to reduce uncertainty in the posterior distribution of a Bayesian inverse analysis. % Key aspects being tested In particular, we examine the relative increase in information gain (RIIG) when moving from single-physics to multi-physics observations. % Description of the test cases Section 4.1 presents a Bayesian inverse problem based on a scalar-valued, one-way coupled electromechanical model. Using this model, we compute the RIIG for varying numbers of observations and noise levels of the second physical field. In Section 4.2, we use a fully-coupled mixed-dimensional multiphase porous medium model to solve a Bayesian inverse problem. Here, we vary the noise levels in the first and second observed fields as well as the coupling strength to assess the impact on the RIIG. % Additional information / Disclaimer These inverse problems and models are not claimed to be relevant real-world applications. They are purely selected to showcase the potential of the multi-physics-enhanced approach to reduce uncertainty in the posterior distribution.
@@
% Purpose of the Demonstration / Why do we choose this model? In this simple demonstration example, the task is to infer the Young’s modulus \$E\$ and Poisson’s ratio \$ν\$ of a material. For this Bayesian inverse problem, the vector of uncertain model parameters is \$x= [E, ν]ˆT\$. We start with a mechanical single-physics model and determine the posterior of the parameters after observing displacements from a tensile test. We then demonstrate the improvement in the posterior’s information content by adding a second, one-way coupled physical field as an additional source of information. This second field consists of electrical measurements, assuming the material of interest is electrically conductive. The electrical field  [^2] By electrical field, we refer here to the general physical field of electricity, encompassing quantities such as voltage, electric current, and electrical resistance. In contrast, electric field denotes the vector field that describes the force per unit charge. In the following, we refer explicitly to the general electrical field, not the electric field. depends on the mechanical field through the displacement, which affects the electrical resistance. For simplicity, we assume that the electrical field does not influence the mechanical field. This one-way coupled electromechanical model was inspired by [39].
@@ 4.1.2 Observations
We generated the synthetic observations based on evaluations of the electromechanical model with the ground-truth parameters \$E_true= 11 $1\text{\,}\mathrm{k}\mathrm{P}\mathrm{a}$\$ and \$ν_true= 0.35\$ at varying force values \$F\$. For a given number of observations \$N_obs, j\$ of the \$j\$-th physical field, these force values were equidistant to each other and the minimal and maximal force values of \$F_min = 0.0 $0\text{\,}\mathrm{N}$\$ and \$F_max = $0.4\text{\,}\mathrm{N}$\$. To obtain the observations, we added zero-mean Gaussian noise to the model outputs (compare Equation 3). The variance of the noise \$σ_jˆ2\$ was computed from the signal-to-noise ratio \$SNR_j\$ of each field \$j\$ according to Equation 12. We generated the noise samples using a Sobol’ sequence [40] to ensure they were evenly distributed and representative of the underlying normal distribution, even for a small number of samples. For the mechanical field, we always used the same displacement observations with a relatively low \$SNR_1 = 50\$ and \$N_obs, 1 = 16\$ observations. For the electrical field, we varied the \$SNR_2\$ and the number of observations \$N_obs, 2\$ to evaluate the relative increase in information gain for different configurations  [^3] As mentioned above, this inverse example is not intended to represent any real, practical application. However, the multi-physics-enhanced BIA in this example can be motivated by the unavailability of cross-sectional measurements during deformation. Only top and bottom measurements are accessible. .
@@ 4.1.3 Probabilistic model & inference method
For the prior distribution of the parameters of interest \$E\$ and \$ν\$, we chose a truncated normal distribution according to Table 2. The truncation bounds \$a\$ and \$b\$ ensure that physically implausible values of \$E\$ and \$ν\$ have a probability density of zero. We further used the Gaussian multi-physics log-likelihood function introduced in Equation 4. The likelihood variance \$σ_jˆ2\$ of each field \$j\$ was always identical to the variance used to sample the observational noise as described in Section 4.1.2. With this prior and likelihood, we evaluated the posterior on a \$100×100\$ grid of \$E\$ and \$ν\$. The grid points in each dimension were evenly spaced with respect to the cumulative distribution function of the prior.
@@ 4.1.4 Comparison of a single-physics and two multi-physics-enhanced Bayesian inverse analyses
Figure 10: Left: The relative increase in information gain (RIIG) becomes larger for a higher number of observations \$N_obs, 2\$ and a higher signal-to-noise ratio \$SNR_2\$ in the second, electrical field. Note that both axes are scaled logarithmically. Right: The multi-physics posteriors for second-field observations with \$N_obs, 2\$ and \$SNR_2\$ and as indicated by the circled numbers in the left plot. Having a higher number of observations \$N_obs, 2\$ and a higher \$SNR_2\$ in the second, electrical field results in a more concentrated posterior.
@@ 4.1.5 Relative increase in information gain for varying second-field observations
% Method 2nd figure Having examined two individual multi-physics posteriors and their information gains in Section 4.1.4, we now investigate the relative increase in information gain (RIIG) for a range of second-field observations. For this investigation, we varied the number \$N_obs, 2\$ and \$SNR_2\$ of the electric current observations and computed the resulting RIIG as introduced in Equation 11. We evaluated the RIIG on a \$50 ×12\$ log-spaced grid of \$N_obs, 2\$ and \$SNR_2\$. The displacement observations were constant throughout the investigation, as already shown in Figure 9.
@@
% Results 2nd figure The left plot in Figure 10 shows the RIIG increasing with the number of observations and the SNR of the second-field observations. Each point in the plot corresponds to a multi-physics-enhanced BIA. This is also illustrated by the right plot in Figure 10, which shows the multi-physics posteriors at specific points in the left plot. The posteriors 1 and 2 were already presented in Figure 9. Combining the high \$SNR_2\$ of posterior 1 with the large number of observations \$N_obs, 2\$ of posterior 2 results in posterior 3. This posterior is highly concentrated and has a RIIG of \$RIIG(y_1, obs, y_2, obs)=3.65\$. Our results demonstrate that observations of a one-way coupled field can substantially reduce the uncertainty in the posterior compared to using single-physics observations alone. This reduction is particularly pronounced when the second-field observations have a high SNR and a high number of observations. However, it should be stressed again that even a few second-field observations can considerably impact the uncertainty if their SNR is sufficiently high. The same holds for observations with a low SNR if a sufficient number of observations is available.
@@ 4.2.1 Forward model
% Model introduction & % Purpose of the Demonstration / Why do we choose this model? We model the lung tissue patch with the mixed-dimensional multiphase porous medium model presented in [41]. The model describes the interaction of solid lung tissue with airways and blood vessels. Larger airways and blood vessels are included in the model as discrete zero-dimensional networks. Smaller respiratory and vascular units and the solid lung tissue are modelled by a homogenized porous medium. In the porous medium, each of the resulting three physical fields (tissue \$t\$, air \$a\$, and blood \$b\$) is assigned a volume fraction \$εˆt, εˆa\$, and \$εˆb\$:
@@
We solve this multi-physics model for the unknown primary variables: the tissue displacement \$dˆt\$, the air pressure \$pˆa\$, and the blood pressure \$pˆb\$. The model is implemented and solved using our open-source multi-physics research code 4C [42]. The governing equations of the model are not essential for this work but are derived in detail in [41]. We discretize these equations in space and time, resulting in a coupled system of nonlinear algebraic equations \$fˆt\$, \$fˆa\$, and \$fˆb\$ for each of the three fields (compare Equation 5):
@@
The three-field porous medium model with tissue, air, and blood is a fully coupled multi-physics model. Figure 11 illustrates this coupling on the right. The additional blood model \$M_2\$ depends on the tissue and air fields through the tissue displacement \$dˆt\$  [^4] For the model parameters selected here, the ratio between the air pressure and the blood pressure \$pˆa / pˆb\$ is always less than 1. Thus, the evolution equation from [41] reduces to Equation 18.:
@@
This evolution equation takes into account the initial blood volume fraction at end-expiration \$εˆb_0\$, the determinant of the deformation gradient \$J\$ with \$X\$ being the material coordinates, and the exponential parameter \$k_J\$. This parameter \$k_J\$ determines the coupling strength from the displacement \$dˆt\$ to the blood volume fraction \$εˆb\$. Please refer to [41] for the coupling mechanism from the blood field back to the tissue and air fields.
@@
None of the blocks in the system matrix are zero, which is characteristic of a fully coupled model. See [41] for more details on the numerical implementation and solution process of the porous medium model.

### v2

@@ 2.1 Coupling in multi-physics models
When moving from a single-physics to a multi-physics-enhanced Bayesian inverse analysis, we need a computational model $\mathcal{M}$ that models each of the observed physical fields. We therefore extend the original single-physics model to a multi-physics model. Such a multi-physics model must account for the coupling between the fields. Regarding inverse analysis, the type of coupling influences the information about the uncertain parameters that observations of each physical field can provide. In this section, we take a closer look at the coupling types in multi-physics models with two fields and their potential to improve an inverse analysis.
@@
Uncoupled One-Way Coupled Fully Coupled
Elastic deformation: $\bm{x}$:1 Young’s modulus $\bm{y}_{1}$: Deflection under force $\bm{y}_{2}$: Vibroacoustic signal Fluid dynamics with smoke: $\bm{x}$:1 Fluid viscosity $\bm{y}_{1}$: Fluid velocity $\bm{y}_{2}$: Smoke density Fluid-structure interaction: $\bm{x}$:1 Fluid parameters $\bm{y}_{1}$: Fluid velocity $\bm{y}_{2}$: Structure deformation
Medical imaging: $\bm{x}$:1 Density field $\bm{y}_{1}$: Ultrasound imaging $\bm{y}_{2}$: Computed tomography (CT) scan Heart contraction [undefq]: $\bm{x}$:1 Electrical state variables $\bm{y}_{1}$: Electrocardiogram $\bm{y}_{2}$: Myocardium displacement Natural convection: $\bm{x}$:1 Thermal parameters $\bm{y}_{1}$: Temperature field $\bm{y}_{2}$: Fluid velocity
Marine electromagnetics [undefk]: $\bm{x}$:1 Seawater & subsurface resistivities $\bm{y}_{1}$: Magnetotellurics $\bm{y}_{2}$: Controlled source electromagnetics Lung ventilation [undefr]: $\bm{x}$:1 Lung stiffness field $\bm{y}_{1}$: Air flow at ventilator $\bm{y}_{2}$: EIT voltages Thermo-structure interaction [undefae]: $\bm{x}$:1 Thermal stress field $\bm{y}_{1}$: Structure deformation $\bm{y}_{2}$: Temperature field
@@
The third coupling type that provides inverse information about the parameters $\bm{x}$ in the second field $\bm{y}_{2}$ is presented in the right diagram of Figure 4. Here, both output fields $\bm{y}_{1}$ and $\bm{y}_{2}$ depend on each other, thus the name fully coupled. An explicit dependency of the second field $\bm{y}_{2}$ on the parameters $\bm{x}$ is optional. Implementing and evaluating fully coupled models is challenging due to the larger system matrices and interdependencies between fields. Depending on the coupling strength, these models often require advanced numerical solution strategies to ensure stability and efficiency (see, e.g., [undefae]).
Information about the coupling type of a multi-physics model is provided by the model’s system matrix. We will briefly address this for multi-physics models with two output fields, although the following concepts also apply to models with an arbitrary number of physical fields [undefae]. The physical fields of interest are typically governed by nonlinear partial differential equations that cannot be solved analytically. Instead, their solutions are approximated at discrete points in space and time. After discretization, a fully coupled model with two physical fields can be formulated as a set of nonlinear algebraic equations:
@@ 3 Information gain in multi-physics-enhanced Bayesian inverse problems
In this work, we measure the uncertainty reduction from the prior to the posterior distribution using the information gain. This measure quantifies the amount of information gained about the parameters $\bm{x}$ by including the observations $\bm{y}_{\text{obs}}$. The information gain is a well-established measure in information theory. It is also widely used in Bayesian optimal experimental design to find an experimental design that maximizes the expected information gain [undefaf, undefa, undefag, undefah, undefai]. In contrast to this use case, we use the information gain here to assess the additional reduction in uncertainty in a Bayesian inverse problem achieved by adding observations of a second physical field.
@@
% Introduce entropy and information gain The information gain is based on the difference in (Shannon) entropy between the prior and the posterior distributions. The entropy \$H(x)\$ quantifies the expected information \$I(x)\$ in a random variable \$x\$ with probability distribution \$p(x)\$ [undefaj]:
@@
This work relies on entropy instead of variance to measure uncertainty. While the variance measures only the spread of a random variable around its mean, the entropy captures the uncertainty across the entire support of its probability distribution. % Moreover, the entropy reflects our intuition of uncertainty as the variance of a random variable with a single mode decreases. If a parameter is highly uncertain, we expect its uncertainty to decrease slowly with decreasing variance. However, if the uncertainty in the parameter is already low, we expect a stronger decrease in uncertainty as the variance decreases further. The entropy exhibits this behavior for a univariate, normally distributed random variable, as shown in Figure 5. % Based on these considerations, we rely on the entropy to compare the uncertainty in the prior and posterior distributions of an inverse problem. In particular, we consider the information gain \$IG(y_obs)\$, which is the relative entropy or Kullback-Leibler divergence between the prior distribution \$p(x)\$ and the posterior distribution \$p(x|y_obs)\$ [undefak]:
@@ 4 Numerical demonstration
% Goal of the numerical demonstration In this section, we demonstrate the potential of multi-physics observations and modeling to reduce uncertainty in the posterior distribution of an orginally single-physics Bayesian inverse analysis. % Key aspects being tested In particular, we examine the relative increase in information gain (RIIG) when moving from single-physics to multi-physics observations. % Description of the test cases Section 4.1 presents a Bayesian inverse problem based on a scalar-valued, one-way coupled electromechanical model. Using this model, we compute the RIIG for varying numbers of observations and noise levels of the second physical field. In Section 4.2, we use a fully-coupled mixed-dimensional multiphase porous medium model to solve a Bayesian inverse problem. Here, we vary the noise levels in the first and second observed fields as well as the coupling strength to assess the impact on the RIIG. % Additional information / Disclaimer These inverse problems and models are not claimed to be relevant real-world applications. They are purely selected to showcase the potential of the multi-physics-enhanced approach to reduce uncertainty in the posterior distribution.
@@
% Purpose of the Demonstration / Why do we choose this model? In this simple demonstration example, the task is to infer the Young’s modulus \$E\$ and Poisson’s ratio \$ν\$ of a material. For this Bayesian inverse problem, the vector of uncertain model parameters is \$x= [E, ν]ˆT\$. We start with a mechanical single-physics model and determine the posterior of the parameters after observing displacements from a tensile test. We then demonstrate the improvement in the posterior’s information content by adding a second, one-way coupled physical field as an additional source of information. This second field consists of electrical measurements, assuming the material of interest is electrically conductive. The electrical field  [^2] By electrical field, we refer here to the general physical field of electricity, encompassing quantities such as voltage, electric current, and electrical resistance. In contrast, electric field denotes the vector field that describes the force per unit charge. In the following, we refer explicitly to the general electrical field, not the electric field. depends on the mechanical field through the displacement, which affects the electrical resistance. For simplicity, we assume that the electrical field does not influence the mechanical field. This one-way coupled electromechanical model was inspired by [undefal].
@@ 4.1.2 Observations
We generated the synthetic observations based on evaluations of the electromechanical model with the ground-truth parameters \$E_true= 11 $1\text{\,}\mathrm{k}\mathrm{P}\mathrm{a}$\$ and \$ν_true= 0.35\$ at varying force values \$F\$. For a given number of observations \$N_obs, j\$ of the \$j\$-th physical field, these force values were equidistant to each other and the minimal and maximal force values of \$F_min = 0.0 $0\text{\,}\mathrm{N}$\$ and \$F_max = $0.4\text{\,}\mathrm{N}$\$. To obtain the observations, we added zero-mean Gaussian noise to the model outputs (compare Equation 3). The variance of the noise \$σ_jˆ2\$ was computed from the signal-to-noise ratio \$SNR_j\$ of each field \$j\$ according to Equation 12. We generated the noise samples using a Sobol’ sequence [undefam] to ensure they were evenly distributed and representative of the underlying normal distribution, even for a small number of samples. For the mechanical field, we always used the same displacement observations with a relatively low \$SNR_1 = 50\$ and \$N_obs, 1 = 16\$ observations. For the electrical field, we varied the \$SNR_2\$ and the number of observations \$N_obs, 2\$ to evaluate the relative increase in information gain for different configurations  [^3] As mentioned above, this inverse example is not intended to represent any real, practical application. However, the multi-physics-enhanced BIA in this example can be motivated by the unavailability of cross-sectional measurements during deformation. Only top and bottom measurements are accessible. .
@@ 4.1.3 Probabilistic model & inference method
For the prior distribution of the parameters of interest \$E\$ and \$ν\$, we chose a truncated normal distribution according to Table 2. The truncation bounds \$a\$ and \$b\$ ensure that physically implausible values of \$E\$ and \$ν\$ have a probability density of zero. Assuming i.i.d. Gaussian noise with known variances, we used the Gaussian multi-physics log-likelihood function introduced in Equation 4. The likelihood variance \$σ_jˆ2\$ of each field \$j\$ was always identical to the variance used to sample the observational noise as described in Section 4.1.2. With this prior and likelihood, we evaluated the posterior on a \$100×100\$ grid of \$E\$ and \$ν\$. The grid points in each dimension were evenly spaced with respect to the cumulative distribution function of the prior. We also repeated the following analyses with finer grids of \$E\$ and \$ν\$ to confirm the overall trends.
@@ 4.1.4 Comparison of a single-physics and two multi-physics-enhanced Bayesian inverse analyses
Figure 10: Top: The relative increase in information gain (RIIG) becomes larger for a higher number of observations \$N_obs, 2\$ and a higher signal-to-noise ratio \$SNR_2\$ in the second, electrical field. Note that both axes are scaled logarithmically. Bottom: The prior \$p(x)\$, the likelihoods of each observed field \$p(y_1, obs|x)\$ and \$p(y_2, obs|x)\$, and the posterior \$p(x|y_1, obs, y_2, obs)\$ for three different multi-physics Bayesian inverse analyses (BIAs). The three BIAs differ in \$N_obs, 2\$ and \$SNR_2\$ of the second-field observations as indicated by the circled numbers in the top plot. The likelihood of the second field increasingly constrains the posterior for higher \$N_obs, 2\$ and \$SNR_2\$.
@@ 4.1.5 Relative increase in information gain for varying second-field observations
% Method 2nd figure Having examined two individual multi-physics posteriors and their respective (relative increase) in information gain in Section 4.1.4, we now investigate the relative increase in information gain (RIIG) for a range of second-field observations. For this investigation, we varied the number \$N_obs, 2\$ and \$SNR_2\$ of the electric current observations and computed the resulting RIIG as introduced in Equation 11. We evaluated the RIIG on a \$50 ×12\$ log-spaced grid of \$N_obs, 2\$ and \$SNR_2\$. The displacement observations were constant throughout the investigation, as already shown in Figure 9.
@@
% Results 2nd figure The top plot in Figure 10 shows the RIIG increasing with the number of observations and the SNR of the second-field observations. Each point in the plot corresponds to a multi-physics-enhanced BIA. This is also illustrated by the bottom plots in Figure 10, which show the prior, the likelihoods of each observed field, and the multi-physics posteriors at specific points in the top plot. The posteriors 1 and 2 were already presented in Figure 9. Although their likelihoods of the two observed fields are widely spread individually, they intersect only in a small region of the parameter space, leading to concentrated posteriors and an RIIG of 1.23 and 1.22, respectively. Combining the high \$SNR_2\$ of posterior 1 with the large number of observations \$N_obs, 2\$ of posterior 2 results in posterior 3. This posterior is highly concentrated and has a RIIG of \$RIIG(y_1, obs, y_2, obs)=3.65\$. Our results demonstrate that observations of a one-way coupled field can substantially reduce the uncertainty in the posterior compared to using single-physics observations alone. This reduction is particularly pronounced when the second-field observations have a high SNR and a high number of observations. However, it should be stressed again that even a few second-field observations can considerably impact the uncertainty if their SNR is sufficiently high. The same holds for observations with a low SNR if a sufficient number of observations is available.
@@ 4.2 Multi-physics-enhanced Bayesian inverse analysis with a complex, fully coupled porous medium model
[anchor=south west,inner sep=0] (img) at (0,0) Refer to caption;
{scope}
@@ 4.2.1 Forward model
% Model introduction & % Purpose of the Demonstration / Why do we choose this model? We model the lung tissue patch with the mixed-dimensional multiphase porous medium model presented in [undefan]. The model describes the interaction of solid lung tissue with airways and blood vessels. Larger airways and blood vessels are included in the model as discrete zero-dimensional networks. Smaller respiratory and vascular units and the solid lung tissue are modelled by a homogenized porous medium. In the porous medium, each of the resulting three physical fields (tissue \$t\$, air \$a\$, and blood \$b\$) is assigned a volume fraction \$εˆt, εˆa\$, and \$εˆb\$:
@@
We solve this multi-physics model for the unknown primary variables: the tissue displacement \$dˆt\$, the air pressure \$pˆa\$, and the blood pressure \$pˆb\$. The model is implemented and solved using our open-source multi-physics research code 4C [undefao]. The governing equations of the model are not essential for this work but are derived in detail in [undefan]. We discretize these equations in space and time, resulting in a coupled system of nonlinear algebraic equations \$fˆt\$, \$fˆa\$, and \$fˆb\$ for each of the three fields (compare Equation 5):
@@
The three-field porous medium model with tissue, air, and blood is a fully coupled multi-physics model. Figure 11 illustrates this coupling on the right. The additional blood model \$M_2\$ depends on the tissue and air fields through the tissue displacement \$dˆt\$  [^4] For the model parameters selected here, the ratio between the air pressure and the blood pressure \$pˆa / pˆb\$ is always less than 1. Thus, the evolution equation from [undefan] reduces to Equation 18.:
@@
This evolution equation takes into account the initial blood volume fraction at end-expiration \$εˆb_0\$, the determinant of the deformation gradient \$J\$ with \$X\$ being the material coordinates, and the exponential parameter \$k_J\$. This parameter \$k_J\$ determines the coupling strength from the displacement \$dˆt\$ to the blood volume fraction \$εˆb\$. Please refer to [undefan] for the coupling mechanism from the blood field back to the tissue and air fields.
@@
None of the blocks in the system matrix are zero, which is characteristic of a fully coupled model. See [undefan] for more details on the numerical implementation and solution process of the porous medium model.